jigglypuff314 (jigglypuff314):

\[If~\int\limits f(x)~sinx~dx~=~-f(x)cosx+\int\limits 3x^2cosx~dx\]then f(x) could be a) 3x^2 b) x^3 c) -x^3 d) sinx e) cosx

jigglypuff314 (jigglypuff314):

Could someone check my work? o-o uv - ∫v du if dv = sinx dx then v = -cosx and then if du = 3x^2 dx then u = 3(1/3)x^3 = x^3 so... B ?

OpenStudy (anonymous):

your question is \[f(x)=\int\limits 3 x^2~dx=\frac{ 3 x^3 }{ 3 }=x^3\] again you are correct.

OpenStudy (anonymous):

yw